OpenStudy (anonymous):

What is the inverse of y = 4x - 2

OpenStudy (anonymous):

A. y = 2x + 4

OpenStudy (anonymous):

\[y= \frac{ 1 }{ 4 }x+\frac{ 1 }{ 2 }\] \[\frac{ 1 }{ 4 }x-\frac{ 1 }{ 2 }\] \[4x+\frac{ 1 }{ }\]

OpenStudy (anonymous):

@ChemicalX Help!! :/

OpenStudy (anonymous):

the last one is supposed to be 4x+1/2

OpenStudy (anonymous):

I think it's A....cuz the inverse is opposite...so yeah. A would be right. :)

OpenStudy (anonymous):

alright sweet. This stuff sucks.. YOu vout to help me mass lol

OpenStudy (anonymous):

lol okay xD

OpenStudy (anonymous):

I"ll tag you in them lol